(y^¹)^2= 1-y^2/1-x^2 ; y=1/2 when, x =1
1
Expert's answer
2021-10-26T16:50:56-0400

Let us solve the differential equation "(y')^2= \frac{1-y^2}{1-x^2} ;" "y=\frac{1}{2}" when "x =1."

This equation is eqivalent to the set of equations "y'= \sqrt{\frac{1-y^2}{1-x^2}}""y'= -\sqrt{\frac{1-y^2}{1-x^2}} ." It follows that "\frac{ dy}{\sqrt{1-y^2} }= \frac{dx}{\sqrt{1-x^2}}" or "\frac{ dy}{\sqrt{1-y^2} }= -\frac{dx}{\sqrt{1-x^2}}". Therefore, "\int\frac{ dy}{\sqrt{1-y^2} }= \int\frac{dx}{\sqrt{1-x^2}}" or "\int\frac{ dy}{\sqrt{1-y^2} }= -\int\frac{dx}{\sqrt{1-x^2}}," and hence "\arcsin y=\arcsin x+C_1" or "\arcsin y=-\arcsin x+C_2." Since "y=\frac{1}{2}" when "x =1," we get that "\arcsin \frac{1}2=\arcsin 1+C_1" or "\arcsin \frac{1}2=-\arcsin 1+C_2."

It follows that "\frac{\pi}6=\frac{\pi}2+C_1,\ \frac{\pi}6=-\frac{\pi}2+C_1," and hence "C_1=-\frac{\pi}{3},\ C_2=\frac{2\pi}{3}."

We conclude that the solutions are the following:

"\arcsin y=\arcsin x-\frac{\pi}{3},\ \arcsin y=-\arcsin x+\frac{2\pi}{3}."